Sunday 10 August 2008
Weather: sunny, occasional cloud.
On site: Mick James, Muriel Hardman, Chris Hobbs, Nigel Harper-Scott, Luke Gearing, Andrew Dear, Greg Ford, Pauline Gimson, Tony Driscoll, Samuel Nobbs, Kyle Brunton, Matthew Dale, Ernie Ford, David Gimson, Anne Pegrum, Andy Nobbs, Philip Dean.
Lorna Boyd-Bell and Rosemary Ryden cried off sick, no contact from Freddy Sharman so not sure why he did not appear.
Trowelling continued on the topsoil, context (1), at the house platform end of the trench. Some excavators started to report occasional chalk flecks associated with a harder feel to the ground. These are, in the main, very small and appear to occur throughout the trench. Not sure if this represents a different context but proceeding on the basis that it is until confirmation (or otherwise) of KJF-M. (Note from KJF-M: it will be best to treat this as a new context).
Trowelling not too bad after the overnight rain initially but the strong wind is drying things out quickly.
Kyle Brunton went home at the start of the afternoon session as his hands were hurting. Matthew Dale went home as he was feeling ill. Both were picked up by their respective mother.
Trowelling became much harder after lunch. Diverted a number of excavators to the hollow way end of the trench as this still had the benefit of last nights rain under an initial crust. Actually looks damp at the bottom of the hollow way.
A number of small pieces of pot coming up which all seem to be associated with holes into an obvious earlier context which is showing in many places through context (1). Substantial number of chalk inclusions. This promises to be a more finds-rich context, I think.
Towards the end of the day, I trowelled a small area of context (1) near the tree stump hollow in the hollow way. Almost immediately came down onto a hard chalk rubble context which extended to the side of the trench. It will be interesting to see if this is the same context as the other holes showing through further up the opposite slope.
Finished the day at 16.45, cleared the site by 17.00